The countries of Utopia and Greatland have the following production possibilities:

Greatland: lumber- 50 Computers- 75

Utopia: lumbar- 80 Computers- 100

a) Which country has absolute advantage in lumber? In computers?

b) Which country has comparative advantage in lumber? In computers?

c) If Utopia produces 40 lumber, how many computers can it produce? If Greatland produces 25 computers, how many lumber can it produce?

d) Fill out the opportunity cost table below:

Greatland 1 lumbar 1 computer

Utopia 1 lumber 1 computer

e) Using separate diagrams, draw Greatland’s PPF and Utopia’s PPF below (put lumber on the upper axis)

f) Assume both countries completely specialize in the good for which they have comparative advantage. On your graphs, show how a trade of 30 lumber for 40 computers will make both countries better off.
Answer #1

a) Absolute advantage means producing a good or service more efficiently than the other country. In this case utopia can produce both lumber and computer more than Greatland and hence has the absolute advantage in production of both the goods. A country can have absolute advantage in both the goods but would have comparative advantage in only one good.

b) Comparative advantage means producing a good or service with lower opportunity cost than other country or good.

In this case for Greatland, it can produce either 50 lumber or 75 computer.

So the opportunity cost to produce one lumber would be 75/50=1.5 computer and the opportunity cost to produce one computer would be 50/75=0.67 lumber

In case of Utopia, the country can produce either 80 lumber or 100 computer. The opportunity cost to produce one lumber would be 100/80=1.25 and the opportunity cost to produce one computer would be 80/100=0.8 lumber.

So Utopia would produce lumber since it has lower opportunity cost and hence has the comparative advantage in that and Greatland would produce computer since the opportunity cost is less for computers in lumber and hence Greatland has comparative advantage in computer.

c) If Utopia produces 40 lumber, then it is giving up 1.25*40= 50 computers ie the opportunity cost to produce 40 lumbers. hence it would be able to produce 50 computers since 50 of them are being given up.

If Greatland produces 25 computers, then the number of lumber given up would be 0.67*25=16.67, and would be able to produce 33.34 lumbers.

d) The opportunity costs were found as above,

Greatland: 1 lumber= 75/50= 1.5 computer

1 computer= 50/75=0.67 lumber

Utopia: 1 lumber= 100/80 1.25 computer

1 computer= 80/100=0.8 lumber